Please format answers in an excel worksheet or make it in an excel worksheet format.

22. [EXCEL] Income statement: Sosa Corporation recently reported an EBITDA of $31.3 million and net income of $9.7 million. The company had $6.8 million in interest expense, and it's average corporate tax rate was 35 percent. What was its depreciation and amortization expense?

23. [EXCEL] Income statement: Fraser Corporation has announced that its net income for the year ended June 30, 2017, was $1,353,412. The company had EBITDA of $4,967,855 and its depreciation and amortization expense was equal to $1,112,685. The company's average tax rate is 34 percent. What was its interest expense?

24. [EXCEL] Income Statement: For its most recent fiscal year, Carmichael Hobby Shop recorded EBITDA of $512,725.20, EBIT of $362,450.20, zero interest expense, and cash flow to investors from operating activity of $348,461.25. Assuming there are no non-cash revenues recorded on the income statement, what is the firm's net income after taxes?
Answer #1

Alignment Number Clipboard B50 Font =B4-B6 F 1 =B4-B6 =B7+B8 2 Answer to Question No. 22 3 Income Statement 4 EBITDA 5 Deprec